# Reinforced Generation of Combinatorial Structures: Applications to Complexity Theory

### v4

@@ Appendix A Proofs of Average Case Theorems
As with previous refutation algorithms [23, 18], our efficient certificate will be a bound on the spectral expansion of a random graph $G\sim G(n,d)$, which is defined as $\lambda^{*}(G)=\max_{i>1}|\lambda_{i}|$ where $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}$ are the eigenvalues of the adjacency matrix of $G$. Friedman’s theorem proves that $\lambda^{*}$ concentrates around $2\sqrt{d-1}$.
Theorem A.1 ([12]).
For all $d\in\mathbb{N}$ $\varepsilon>0$, with probability $1-o_{n}(1)$ over $G\sim G(n,d)$, we have $\lambda^{*}(G)\leq 2\sqrt{d-1}+\varepsilon$.
Definition A.1 ($d$-ary tree).
@@
Our improvement on Hoffman’s classical bounds consists of concluding stronger bounds on $\textup{MC}(G)$ and $\textup{IS}(G)$ given that $\lambda^{*}(G)\leq\lambda$ for $\lambda=2\sqrt{d-1}+\varepsilon$. Let us denote by $\Omega_{d,\lambda}$ the set of $d$-regular graphs $G$ such that $\lambda^{*}(G)\leq\lambda$. We will require the notion of a labeled $d$-ary tree, which we define below.
Definition A.2 (Labeled $d$-ary tree).
@@
Let $G\in\Omega_{d,\lambda}$ be an $n$-vertex graph, and let $S\subseteq V(G)$ be a cut in $G$. Let $\alpha\in[0,1/2]$ be a value that approximates $|S|/n$, in the sense that $|\alpha-|S|/n|\leq\delta$. Define the cut vector $x\in\{-1,1\}^{n}$ by $x_{i}=(-1)^{\mathbf{1}\{i\in S\}}$. We will associate the pair $(G,S)$ with a distribution $\mu_{G,S}\in\Lambda_{d,L}$, which can be sampled from as follows:
@@
Lemma A.1 (Local Consistency Constraints).
@@
Lemma A.2 (Average Value of root).
@@
Lemma A.3 (Spectral Constraints).
@@
$\mathbb{E}_{y\sim\mu_{G,S},\ell}[y_{0}\cdot y_{\ell}]=\frac{x^{T}(A/d)^{L}x}{n}=\frac{x^{T}A^{L}x}{d^{L}\cdot n}.$
Let us write $A=d\frac{\mathbf{1}\mathbf{1}^{T}}{n}+\widetilde{A}$, where $d=\lambda_{1}$ is the trivial eigenvalue of $A$ and $\widetilde{A}$ is the portion of $A$ orthogonal to the trivial eigenvector $\mathbf{1}$. That is, $\widetilde{A}\mathbf{1}=\mathbf{0}$. Note that $\|\widetilde{A}\|_{op}\leq\lambda$. Let us proceed with our computation.
$\mathbb{E}_{y\sim\mu_{G,S},\ell}[y_{0}\cdot y_{\ell}]$ $=\frac{d^{L}\cdot x^{T}\mathbf{1}\mathbf{1}^{T}x}{d^{L}\cdot n^{2}}+\frac{x^{T}\widetilde{A}^{L}x}{d^{L}\cdot n}$
$=\frac{(2|S|-n)^{2}}{n^{2}}+\frac{x^{T}\widetilde{A}^{L}x}{d^{L}\cdot n}.$
The Lemma follows from (1) the fact that $|\alpha-|S|/n|\leq\delta$, and (2) the bounds $|x^{T}\widetilde{A}^{L}x|\leq\|\widetilde{A}\|_{op}\cdot\|x\|_{2}^{2}\leq\lambda n$, along with the bound $x^{T}\widetilde{A}^{L}x\geq 0$ if $L$ is even. ∎
Lemma A.4 (Objective Value).
@@
$\mathbb{E}_{y\sim\mu_{G,S},i\sim[d]}[y_{0}\cdot y_{i}]=\frac{x^{T}Ax}{nd}=1-2\frac{|\delta_{G}(S,\bar{S})|}{|E(G)|},$
@@
Lemmas A.1, A.2 and A.3 are all linear constraints in the pdf of the distribution $\mu=\mu_{G,S}$, and Lemma A.4 is a linear objective function. We may solve this LP in the variables $\mu$ for small values of $L$ and $d$ to obtain valid upper bounds on objective function, the cut value $|\delta_{G}(S,\bar{S})|/|E(G)|$. If we solve this LP for a collection of discrete values $\alpha\in\{0,\delta,2\delta,\ldots,1\}$, we obtain an upper bound on the cut value of any cut of a graph $G\in\Omega_{d,\lambda}$.
@@
We remark that the above LP appears intractable – in particular, it has $2^{46}$ variables even for $d=3$, $L=4$. We apply two optimizations to make it more tractable in practice.
@@
    Use the same variable for isomorphic labelings of $T_{d,L}$. In particular, if there is a graph isomorphism of $T_{d,L}$ that maps a labeling $y$ to another labeling $y^{\prime}$, then Lemma A.1 implies that they should receive the same probability in $\mu$.
@@
    Only consider “locally optimal” labelings of $T_{d,L}$. To do this, we apply a preprocessing step, where we ensure that the cut $S$ is locally optimal in the sense that for any vertex $i\in S$, the labeling of $T_{d,L}$ generated by restricting $x\in\{\pm 1\}^{n}$ to the $L$-step neighborhood of $i$ corresponds to a maximum cut of $T_{d,L}$, subject to the “boundary conditions” of labels of the leaves of $T_{d,L}$. Therefore, we can also discard any variables corresponding to such suboptimal labelings of $T_{d,L}$.
@@
Appendix B Proofs of Theorems 4.1 and 4.2
In this section, we give the details of the NP-hardness of approximation proofs for MAX-k-CUT for $k\in\{3,4\}$. Let us begin by formally defining a gadget, specifically in the context of the reduction from $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$. We will need $k$ separate gadgets. In particular, for each defining predicate $P_{i}^{=}$ of $3\textup{LIN}({k})$, we will need an $i$-gadget, which we define below.
Definition B.1 (Gadgets).
Let $k\geq 2$ and define $\Sigma=\mathbb{Z}_{k}$.
  • $i$-gadget: For $i\in\Sigma$, an $i$-gadget from $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$ is an instance $I_{i}$ of $\textup{MAX-}{k}\textup{-CUT}$ over $3+k+n_{\textup{aux}}$ variables, where $n_{\textup{aux}}\in\mathbb{N}$.
  • Primary, global, and auxiliary variables: We partition the set of variables $3+k+n_{\textup{aux}}]$ into three primary variables $\{1,2,3\}$, $k$ global variables $\{4,5,\ldots,k+3\}$, and $n_{\textup{aux}}$ auxiliary variables $\{k+4,\ldots,3+k+n_{\textup{aux}}\}$. We will always write an assignment to the variables as a tuple $(x,y,z)\in\Sigma^{3+k+n_{\textup{aux}}}$ where $x\in\Sigma^{3}$, $y\in\Sigma^{k}$, and $z\in\Sigma^{n_{\textup{aux}}}$.
The variables $y$ will be restricted in a way to break the symmetries of $\textup{MAX-}{k}\textup{-CUT}$. Concretely, let $Y$ be the set of strings $y$ in $\Sigma^{k}$ that are lexicographically not larger than any string $y^{\sigma}=(\sigma(y_{1}),\sigma(y_{2}),\ldots,\sigma(y_{k}))$ that is obtained by permuting the alphabet $\Sigma$ by some permutation $\sigma:\Sigma\to\Sigma$. In particular, note that the vector $(0,1,2,\ldots,k-1)$ is in $Y$. We will only ever assign $y$ to values in $Y$.
Definition B.2 (Soundness and completeness).
@@
We will associate four real-valued parameters to the gadget $I_{i}=\sum_{j\in[m]}c_{j}$ as follows.
  • (Completeness analysis) $c(I_{i})$ is defined as $\min_{x\in C_{i}}\max_{z\in\Sigma^{n_{\textup{aux}}}}I_{i}(x,y,z)$ for $y=(0,1,\ldots,k-1)$. For $x\in C_{i}$, we say that $\arg\max_{z\in\Sigma^{n_{\textup{aux}}}}I_{i}(x,y,z)$ is the witness for $x$.
  • (Soundness analysis) $c^{\prime}(I_{i})$ is defined as
    $c^{\prime}(I_{i})=\max_{x\in C_{i},z\in\Sigma^{n_{\textup{aux}}},y\in Y}I_{i}(x,y,z).$
    Also, $s(I_{i})$ is defined as
    $s(I_{i})=\max_{x\in S_{i},z\in\Sigma^{n_{\textup{aux}}},y\in Y}I_{i}(x,y,z).$
@@
We will show that any collection of $i$-gadgets for all $i\in\mathbb{Z}_{k}$ provides a reduction whose performance depends on the various parameters defined above. After this, in Section B.1, we give an explicit description of our gadgets and state the parameters achieved by them.
@@
Theorem B.1 ([19]).
Let $k\geq 2$. Given an instance $I=\sum_{j\in[m]}c_{j}$ of $3\textup{LIN}({k})$ with $m/k$ clauses from each of the predicates $P_{0}^{=},P_{1}^{=},\ldots,P_{k-1}^{=}$, it is NP-hard to distinguish between the following cases for any $\varepsilon>0$:
@@
    (Completeness case) $\max_{x}I(x)\geq(1-\varepsilon)\cdot m$, or
@@
    (Soundness case) $\max_{x}I(x)\leq(1/k+\varepsilon)\cdot m$.
Theorem B.2 details how our gadgets imply NP-hardness of approximation for $\textup{MAX-}{k}\textup{-CUT}$, with $k\in\{3,4\}$.
Theorem B.2.
Let $k\geq 2$, and set $\Sigma=\mathbb{Z}_{k}$. Let $I_{i}$ be an $i$-gadget from $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$ for each $i\in\Sigma$. For any $\varepsilon>0$, given an instance $J$ of $\textup{MAX-}{k}\textup{-CUT}$ with $M$ constraints, it is NP-hard to distinguish between the following two cases for any $\varepsilon>0$:
@@
    (Completeness case) $\max_{x}J(x)\geq(a-\varepsilon)\cdot M$, or
@@
    (Soundness case) $\max_{x}J(x)\leq(b+\varepsilon)\cdot M$,
@@
$a=\frac{\sum_{i\in\Sigma}c(I_{i})}{\sum_{i\in\Sigma}t(I_{i})},\quad b=\frac{c^{\prime}(I_{r_{0}})+\sum_{i\in\Sigma\setminus\{0\}}s(I_{r_{i}})}{\sum_{i\in\Sigma}t(I_{i})}.$
Here $r_{0},r_{1},\ldots r_{k-1}$ is a permutation of $\{0,1,\ldots,k-1\}$ such that $c^{\prime}(I_{r_{i}})-s(I_{r_{i}})$ is maximized by $r_{0}$.
Proof of Theorem B.2.
We will perform a reduction starting from the NP-hard problem in Theorem B.1 to $\textup{MAX-}{k}\textup{-CUT}$. Let $I=\sum_{j\in[m]}c_{j}$ be an $n$-variable instance of $3\textup{LIN}({k})$, and suppose that it has $m/k$ clauses using $P_{i}^{=}$ for all $i\in\Sigma$. We will show how to reduce $I$ to an instance $J$ of $\textup{MAX-}{k}\textup{-CUT}$.
Let $n_{\textup{aux}}$ be the number of auxiliary variables in all gadgets $I_{i}$; we can assume they all have the same number of auxiliary variables by introducing new dummy variables to some of them. $J$ will be defined as an instance on $n+k+n_{\textup{aux}}\cdot m$ variables, which we will denote as $x\in\Sigma^{n}$, $y\in\Sigma^{k}$, and $z\in\Sigma^{n_{\textup{aux}}\cdot m}$. For $j\in[m]$ we will denote the $j^{th}$ block of size $n_{\textup{aux}}$ in $z$ as $z^{j}$. We will refer to $x,y,z$ as the primary, global, and auxiliary variables respectively.
Let $c_{j}=P^{=}_{i}(x_{j_{1}},x_{j_{2}},x_{j_{3}})$ be a clause corresponding to the predicate $P^{=}_{i}$ for some $i$. We will add a copy of $I_{i}$ on the variables $(x^{\prime},y,z^{j})$, where $x^{\prime}=(x_{j_{1}},x_{j_{2}},x_{j_{3}})$. Note that the number of clauses in $J$ is $M=(m/k)\cdot\sum_{i\in\Sigma}t(I_{i})$.
Now we will invoke Theorem B.1, which says that it is NP-hard to distinguish between the cases that $I$ is almost completely satisfiable, and $I$ is essentially $1/k$-satisfiable.
@@
    (Completeness case). Assuming there is an $x\in\Sigma^{n}$ such that $I(x)\geq(1-\varepsilon)\cdot m$, we show how to construct an assignment $(x,y,z)\in\Sigma^{n+k+n_{\textup{aux}}\cdot m}$ such that $J(x,y,z)\geq a\cdot M$. We will set $y=(0,1,2,\ldots,k-1)$, and for each clause $c_{j}=P_{i}^{=}(x_{j_{1}},x_{j_{2}},x_{j_{3}})$, $z^{j}$ will be set as the witness $z_{j}=\arg\max_{z^{j}\in\Sigma^{n_{\textup{aux}}}}I_{i}(x^{\prime},y,z^{j})$, where $x^{\prime}=(x_{j_{1}},x_{j_{2}},x_{j_{3}})$.
    It is immediate that $(x,y,z)$ satisfies at least
    $\sum_{i\in\Sigma}c(I_{i})\cdot m/k-O(\varepsilon\cdot m)=a\cdot M-O(\varepsilon)\cdot m\geq(a-\varepsilon^{\prime})\cdot M$
@@
    (Soundness case). We must show that if $\max_{x}I(x)\leq(1/3+\varepsilon)\cdot m$, then $\max_{x,y,z}J(x,y,z)\leq(b+\varepsilon^{\prime})\cdot M$. Let $(x,y,z)$ be an assignment to the variables of $J$.
    The first step we perform is to assume without loss of generality that $y\in Y$. To do this, note that by definition of $Y$, there is some permutation $\sigma:\Sigma\to\Sigma$ such that $y^{\sigma}=(\sigma(y_{1}),\ldots,\sigma(y_{k}))\in Y$. We will apply $\sigma$ to all variables in the full assignment $(x,y,z)$, resulting in no change in the value of $J(x,y,z)$ while guaranteeing that $y\in Y$.
    For $i\in\Sigma$, let $\alpha_{i}$ denote the fraction of $P^{=}_{i}$ constraints of $I$ that are satisfied by $x$. By assumption, we have $\sum_{i\in\Sigma}\alpha_{i}\leq 1+3\varepsilon$. Let us define $\hat{\alpha_{i}}=\alpha_{i}/(\sum_{i^{\prime}\in\Sigma}\alpha_{i^{\prime}})$, so $\sum_{i\in\Sigma}\hat{\alpha_{i}}=1$. In other words, $\{\hat{\alpha_{i}}:i\in\Sigma\}$ defines a probability distribution.

### v5

@@ Appendix B Proofs of Average Case Theorems
As with previous refutation algorithms [30, 25], our efficient certificate will be a bound on the spectral expansion of a random graph $G\sim\mathcal{G}(n,d)$, which is defined as $\lambda^{*}(G)=\max_{i>1}|\lambda_{i}|$ where $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}$ are the eigenvalues of the adjacency matrix $\mathbf{A}$ of the graph $G$. Friedman’s theorem proves that $\lambda^{*}$ concentrates around $2\sqrt{d-1}$.
Theorem B.1 ([18]).
For all $d\in\mathbb{N}$ $\varepsilon>0$, with probability $1-o_{n}(1)$ over $G\sim\mathcal{G}(n,d)$, we have $\lambda^{*}(G)\leq 2\sqrt{d-1}+\varepsilon$.
Definition B.1 ($d$-ary tree).
@@
Our improvement on Hoffman’s classical bounds [30, 25] consists of concluding stronger bounds on $\textup{MC}(G)$ and $\textup{IS}(G)$ given that $\lambda^{*}(G)\leq\lambda$ for $\lambda=2\sqrt{d-1}+\varepsilon$. Let us denote by $\Omega_{d,\lambda}$ the set of $d$-regular graphs $G$ such that $\lambda^{*}(G)\leq\lambda$. We will require the notion of a labeled $d$-ary tree, which we define below.
Definition B.2 (Labeled $d$-ary tree).
@@
Let $G\in\Omega_{d,\lambda}$ be an $n$-vertex graph, and let $S\subseteq V(G)$ be a cut in $G$. Let $\alpha\in[0,1/2]$ be a value that approximates $|S|/n$, in the sense that $|\alpha-|S|/n|\leq\delta$. Define the cut vector $\mathbf{x}\in\{-1,1\}^{n}$ by $x_{i}=(-1)^{\mathbf{1}\{i\in S\}}$. We will associate the pair $(G,S)$ with a distribution $\mu_{G,S}\in\Lambda_{d,L}$, which can be sampled from as follows:
@@
Lemma B.1 (Local Consistency Constraints).
@@
Lemma B.2 (Average Value of root).
@@
Lemma B.3 (Spectral Constraints).
@@
$\mathbb{E}_{y\sim\mu_{G,S},\ell}[y_{0}\cdot y_{\ell}]=\frac{\mathbf{x}^{\top}(\mathbf{A}/d)^{L}\mathbf{x}}{n}=\frac{\mathbf{x}^{\top}\mathbf{A}^{L}\mathbf{x}}{d^{L}\cdot n}.$
Let us write $\mathbf{A}=d\frac{\mathbf{1}\mathbf{1}^{\top}}{n}+\widetilde{\mathbf{A}}$, where $d=\lambda_{1}$ is the trivial eigenvalue of $\mathbf{A}$ and $\widetilde{\mathbf{A}}$ is the portion of $\mathbf{A}$ orthogonal to the trivial eigenvector $\mathbf{1}$. That is, $\widetilde{A}\cdot\mathbf{1}=\mathbf{0}$. Note that $\|\widetilde{\mathbf{A}}\|_{\sf op}\leq\lambda$. Let us proceed with our computation.
$\mathbb{E}_{y\sim\mu_{G,S},\ell}[y_{0}\cdot y_{\ell}]$ $=\frac{d^{L}\cdot\mathbf{x}^{\top}\mathbf{1}\mathbf{1}^{T}\mathbf{x}}{d^{L}\cdot n^{2}}+\frac{\mathbf{x}^{\top}\widetilde{\mathbf{A}}^{L}\mathbf{x}}{d^{L}\cdot n}$
$=\frac{(2|S|-n)^{2}}{n^{2}}+\frac{\mathbf{x}^{\top}\widetilde{\mathbf{A}}^{L}\mathbf{x}}{d^{L}\cdot n}.$
The Lemma follows from (1) the fact that $|\alpha-|S|/n|\leq\delta$, and (2) the bounds $|\mathbf{x}^{\top}\widetilde{\mathbf{A}}^{L}\mathbf{x}|\leq\|\widetilde{\mathbf{A}}\|_{\sf op}\cdot\|\mathbf{x}\|_{2}^{2}\leq\lambda n$, along with the bound $\mathbf{x}^{\top}\widetilde{A}^{L}\mathbf{x}\geq 0$ if $L$ is even. ∎
Lemma B.4 (Objective Value).
@@
$\mathbb{E}_{y\sim\mu_{G,S},i\sim[d]}[y_{0}\cdot y_{i}]=\frac{\mathbf{x}^{\top}\mathbf{A}\mathbf{x}}{nd}=1-2\frac{|\delta_{G}(S,\bar{S})|}{|E(G)|},$
@@
Lemmas B.1, B.2 and B.3 are all linear constraints in the pdf of the distribution $\mu=\mu_{G,S}$, and Lemma B.4 is a linear objective function. We may solve this LP in the variables $\mu$ for small values of $L$ and $d$ to obtain valid upper bounds on objective function, the cut value $|\delta_{G}(S,\bar{S})|/|E(G)|$. If we solve this LP for a collection of discrete values $\alpha\in\{0,\delta,2\delta,\ldots,1\}$, we obtain an upper bound on the cut value of any cut of a graph $G\in\Omega_{d,\lambda}$.
@@
The above LP appears intractable – it has $2^{46}$ variables even for $d=3$, $L=4$. We apply two optimizations to make it more tractable in practice.
@@
    Use the same variable for isomorphic labelings of $T_{d,L}$. In particular, if there is a graph isomorphism of $T_{d,L}$ that maps a labeling $y$ to another labeling $y^{\prime}$, then Lemma B.1 implies that they should receive the same probability in $\mu$.
@@
    Only consider “locally optimal” labelings of $T_{d,L}$. To do this, we apply a preprocessing step, where we ensure that the cut $S$ is locally optimal in the sense that for any vertex $i\in S$, the labeling of $T_{d,L}$ generated by restricting $\mathbf{x}\in\{\pm 1\}^{n}$ to the $L$-step neighborhood of $i$ corresponds to a maximum cut of $T_{d,L}$, subject to the “boundary conditions” of labels of the leaves of $T_{d,L}$. Therefore, we can also discard any variables corresponding to such suboptimal labelings of $T_{d,L}$.
@@
Appendix C Proofs of Theorems 3.1 and 3.2
In this section, we give the details of the NP-hardness of approximation proofs for MAX-k-CUT for $k\in\{3,4\}$. For ease of exposition, we begin by reintroducing all predicates needed for the proof. All predicates will be over the alphabet $\Psi=\mathbb{Z}_{k}$.
  • $P^{\neq}$ denotes a $2$-ary predicate defined by $P^{\neq}(x,y)=\mathbf{1}\{x\neq y\}$.
  • For $i\in\{0,\ldots,k-1\}$, $P^{=}_{i}$ denotes a $3$-ary predicate defined by $P^{=}_{i}(x,y,z)=\mathbf{1}\{x+y+z\equiv i\pmod{k}\}$.
[…]
Definition C.1 (Gadgets).
Let $k\geq 2$ and define $\Psi=\mathbb{Z}_{k}$.
  • $i$-gadget: For $i\in\Psi$, an $i$-gadget from $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$ is an instance $I_{i}$ of $\textup{MAX-}{k}\textup{-CUT}$ over $3+k+n_{\textup{aux}}$ variables, where $n_{\textup{aux}}\in\mathbb{N}$.
  • Primary, global, and auxiliary variables: We partition the set of variables $3+k+n_{\textup{aux}}]$ into three primary variables $\{1,2,3\}$, $k$ global variables $\{4,5,\ldots,k+3\}$, and $n_{\textup{aux}}$ auxiliary variables $\{k+4,\ldots,3+k+n_{\textup{aux}}\}$. We will always write an assignment to the variables as a tuple $(\mathbf{x},\mathbf{y},\mathbf{z})\in\Psi^{3+k+n_{\textup{aux}}}$ where $\mathbf{x}\in\Psi^{3}$, $\mathbf{y}\in\Psi^{k}$, and $\mathbf{z}\in\Psi^{n_{\textup{aux}}}$.
The variables $\mathbf{y}$ will be restricted in a way to break the symmetries of $\textup{MAX-}{k}\textup{-CUT}$. Concretely, let $Y$ be the set of strings $\mathbf{y}$ in $\Psi^{k}$ that are lexicographically not larger than any string $\mathbf{y}^{\sigma}=(\sigma(y_{1}),\sigma(y_{2}),\ldots,\sigma(y_{k}))$ that is obtained by permuting the alphabet $\Psi$ by some permutation $\sigma:\Psi\to\Psi$. In particular, note that the vector $(0,1,2,\ldots,k-1)$ is in $Y$. We will only ever assign $\mathbf{y}$ to values in $Y$.
Definition C.2 (Soundness and completeness).
@@
We will associate four real-valued parameters to the gadget $I_{i}=\sum_{j\in[m]}\zeta_{j}$ as follows.
  • (Completeness analysis) $c(I_{i})$ is defined as
    $c(I_{i})=\min_{\mathbf{x}\in C_{i}}\max_{\mathbf{z}\in\Psi^{n_{\textup{aux}}}}I_{i}(\mathbf{x},\mathbf{y},\mathbf{z})$
    for $\mathbf{y}=(0,1,\ldots,k-1)$. For $\mathbf{x}\in C_{i}$, we say that $\arg\max_{\mathbf{z}\in\Psi^{n_{\textup{aux}}}}I_{i}(\mathbf{x},\mathbf{y},\mathbf{z})$ is the witness for $\mathbf{x}$.
  • (Soundness analysis) $s_{1}(I_{i})$ is defined as
    $s_{1}(I_{i})=\max_{\mathbf{x}\in C_{i},\mathbf{z}\in\Psi^{n_{\textup{aux}}},\mathbf{y}\in Y}I_{i}(\mathbf{x},\mathbf{y},\mathbf{z}).$
    Also, $s_{2}(I_{i})$ is defined as
    $s_{2}(I_{i})=\max_{\mathbf{x}\in S_{i},\mathbf{z}\in\Psi^{n_{\textup{aux}}},\mathbf{y}\in Y}I_{i}(\mathbf{x},\mathbf{y},\mathbf{z}).$
@@
We will show that any collection of $i$-gadgets for all $i\in\mathbb{Z}_{k}$ provides a reduction whose performance depends on the various parameters defined above. After this, in Appendix C.1, we give an explicit description of our gadgets and state the parameters achieved by them.
@@
Theorem C.1 ([26]).
Let $k\geq 2$. Given an instance $I=\sum_{j\in[m]}\zeta_{j}$ of $3\textup{LIN}({k})$ with $m/k$ clauses from each of the predicates $P_{0}^{=},P_{1}^{=},\ldots,P_{k-1}^{=}$, it is NP-hard to distinguish between the following cases for any $\varepsilon>0$:
@@
    Completeness case. $\max_{\mathbf{x}}I(\mathbf{x})\geq(1-\varepsilon)\cdot m$, or
@@
    Soundness case. $\max_{\mathbf{x}}I(\mathbf{x})\leq(1/k+\varepsilon)\cdot m$.
Theorem C.2 details how our gadgets imply NP-hardness of approximation for $\textup{MAX-}{k}\textup{-CUT}$, with $k\in\{3,4\}$.
Theorem C.2.
Let $k\geq 2$, and set $\Psi=\mathbb{Z}_{k}$. Let $I_{i}$ be an $i$-gadget from $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$ for each $i\in\Psi$. For any $\varepsilon>0$, given an instance $J$ of $\textup{MAX-}{k}\textup{-CUT}$ with $M$ constraints, it is NP-hard to distinguish between the following two cases for any $\varepsilon>0$:
@@
    Completeness case: $\max_{\mathbf{x}}J(\mathbf{x})\geq(a-\varepsilon)\cdot M$, or
@@
    Soundness case: $\max_{\mathbf{x}}J(\mathbf{x})\leq(b+\varepsilon)\cdot M$,
@@
$a=\frac{\sum_{i\in\Psi}c(I_{i})}{\sum_{i\in\Psi}t(I_{i})},\quad b=\frac{s_{1}(I_{r_{0}})+\sum_{i\in\Psi\setminus\{0\}}s_{2}(I_{r_{i}})}{\sum_{i\in\Psi}t(I_{i})}.$
Here $r_{0},r_{1},\ldots r_{k-1}$ is a permutation of $\{0,1,\ldots,k-1\}$ such that $s_{1}(I_{r_{i}})-s_{2}(I_{r_{i}})$ is maximized by $r_{0}$.
Proof of Theorem C.2.
We will perform a reduction starting from the NP-hard problem in Theorem C.1 to $\textup{MAX-}{k}\textup{-CUT}$. Let $I=\sum_{j\in[m]}\zeta_{j}$ be an $n$-variable instance of $3\textup{LIN}({k})$, and suppose that it has $m/k$ clauses using $P_{i}^{=}$ for all $i\in\Psi$. We will show how to reduce $I$ to an instance $J$ of $\textup{MAX-}{k}\textup{-CUT}$.
Let $n_{\textup{aux}}$ be the number of auxiliary variables in all gadgets $I_{i}$; we can assume they all have the same number of auxiliary variables by introducing new dummy variables to some of them. $J$ will be defined as an instance on $n+k+n_{\textup{aux}}\cdot m$ variables, which we will denote as $\mathbf{x}\in\Psi^{n}$, $\mathbf{y}\in\Psi^{k}$, and $\mathbf{z}\in\Psi^{n_{\textup{aux}}\cdot m}$. For $j\in[m]$ we will denote the $j^{th}$ block of size $n_{\textup{aux}}$ in $z$ as $z^{j}$. We will refer to $\mathbf{x},\mathbf{y},\mathbf{z}$ as the primary, global, and auxiliary variables respectively.
Let $\zeta_{j}=P^{=}_{i}(x_{j_{1}},x_{j_{2}},x_{j_{3}})$ be a clause corresponding to the predicate $P^{=}_{i}$ for some $i$. We will add a copy of $I_{i}$ on the variables $(\mathbf{x}^{\prime},\mathbf{y},\mathbf{z}^{j})$, where $\mathbf{x}^{\prime}=(x_{j_{1}},x_{j_{2}},x_{j_{3}})$. Note that the number of clauses in $J$ is $M=(m/k)\cdot\sum_{i\in\Psi}t(I_{i})$.
Now we will invoke Theorem C.1, which says that it is NP-hard to distinguish between the cases that $I$ is almost completely satisfiable, and $I$ is essentially $1/k$-satisfiable.
@@
    Completeness case: Assuming there is an $\mathbf{x}\in\Psi^{n}$ such that $I(\mathbf{x})\geq(1-\varepsilon)\cdot m$, we show how to construct an assignment $(\mathbf{x},\mathbf{y},\mathbf{z})\in\Psi^{n+k+n_{\textup{aux}}\cdot m}$ such that $J(\mathbf{x},\mathbf{y},\mathbf{z})\geq a\cdot M$. We will set $\mathbf{y}=(0,1,2,\ldots,k-1)$, and for each clause $\zeta_{j}=P_{i}^{=}(x_{j_{1}},x_{j_{2}},x_{j_{3}})$, $\mathbf{z}^{j}$ will be set as the witness $\mathbf{z}^{j}=\arg\max_{\mathbf{z}\in\Psi^{n_{\textup{aux}}}}I_{i}(\mathbf{x}^{\prime},\mathbf{y},\mathbf{z})$, where $\mathbf{x}^{\prime}=(x_{j_{1}},x_{j_{2}},x_{j_{3}})$.
    It is immediate that $(\mathbf{x},\mathbf{y},\mathbf{z})$ satisfies at least
    $\sum_{i\in\Psi}c(I_{i})\cdot m/k-O(\varepsilon\cdot m)=a\cdot M-O(\varepsilon)\cdot m\geq(a-\varepsilon^{\prime})\cdot M$
@@
    Soundness case: We must show that if $\max_{\mathbf{x}}I(\mathbf{x})\leq(1/k+\varepsilon)\cdot m$, then $\max_{\mathbf{x},\mathbf{y},\mathbf{z}}J(\mathbf{x},\mathbf{y},\mathbf{z})\leq(b+\varepsilon^{\prime})\cdot M$. Let $(\mathbf{x},\mathbf{y},\mathbf{z})$ be an assignment to the variables of $J$.
    The first step we perform is to assume without loss of generality that $\mathbf{y}\in Y$. To do this, note that by definition of $Y$, there is some permutation $\sigma:\Psi\to\Psi$ such that $\mathbf{y}^{\sigma}=(\sigma(y_{1}),\ldots,\sigma(y_{k}))\in Y$. We will apply $\sigma$ to all variables in the full assignment $(\mathbf{x},\mathbf{y},\mathbf{z})$, resulting in no change in the value of $J(\mathbf{x},\mathbf{y},\mathbf{z})$ while guaranteeing that $\mathbf{y}\in Y$.
    For $i\in\Psi$, let $\alpha_{i}$ denote the fraction of $P^{=}_{i}$ constraints of $I$ that are satisfied by $\mathbf{x}$. By assumption, we have $\sum_{i\in\Psi}\alpha_{i}\leq 1+k\varepsilon$. Let us define $\hat{\alpha_{i}}=\alpha_{i}/(\sum_{i^{\prime}\in\Psi}\alpha_{i^{\prime}})$, so $\sum_{i\in\Psi}\hat{\alpha_{i}}=1$. In other words, $\{\hat{\alpha_{i}}:i\in\Psi\}$ defines a probability distribution.